A ball is dropped from the top of a 5.00-m high building. If air resistance is neglected, what is its velocity after 0.70 s?




1
Expert's answer
2021-11-16T19:22:06-0500

Initial velocity, "u=0m\/s, t=0.7s"

"v=u+gt=0+9.8\times 0.7=6.86m/s"